OpenStudy (anonymous):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

Nnesha (nnesha):

do you know the formula to find the slope

OpenStudy (anonymous):

A. 0

OpenStudy (triciaal):

the slope is the change in y / corresponding change in x points (x1, y1) and (x2, y2) slope = (y2 - y1)/(x2 - x1) you have 2 points plug your numbers and do the calculations

OpenStudy (anonymous):

D. undefined

OpenStudy (anonymous):

I think It's D.

OpenStudy (triciaal):

you cannot divide by 0 so its undefined

OpenStudy (triciaal):

9 --3/2-2 = 12/0 = undefined yes D